Timeline of instrument status

The following table describes the events affecting the quality or the measurement of data obtained with KMOS. Its content can be downloaded as a Comma Separated Values file. It is complete as of September 13th, 2020.

The Start Date column indicates when the event took place. Some events cannot be defined by a single instant, but extend over sometimes several days, weeks or months. Operation with the instrument may continue while the event occurs, in which case the Ongoing flag is set. If the event had ended, the Ongoing ? flag is lowered and End Date is given. The Type column indicates the characteristic(s) of the event: Earthquake, Maintenance, Upgrade, Element failure, Intervention, Pipeline installation at QC, Change of nominal values or thresholds, Modification of configuration file of this database, or Other. The Instrument Modes column indicates the affected mode. The column KPI indicates which quality control parameters identified as Key Performance Indicators (i.e., parameters following the most critical quantities) are affected; the column non-KPI indicates which other critical Quality Control parameters are affected. The Calibration Raw Types column indicates the raw types of the affected calibrations. The Calibration Breakpoint flag indicates if calibrations obtained before the event cannot be used for science data obtained after the event or if calibration data obtained after the event cannot be used with science data obtained before the event. The Comment column provides more details. The History column allows one to see possible past information on the entry.

The Quality Control pages for KMOS can be accessed through the QC overview page


  • 2022-11-22: since ARM11 (previous bullet) was put in simulation the pipeline has struggled to properly reduce flats and waves for detector 2. The raw data for the arms on detector 2 not in simulation is considered fine, however the pipeline cannot currently reduce these data. For now we are continuing to observe with KMOS and a pipeline fix will be implemeted asap so that data from all arms within detector 2 not in simulation can indeed be reduced.
  • 2022-11-06: ARM11 was found to have problems moving in the angular direction and it was concluded that it had to be put in simulation. Therefore, from 06-11-22 KMOS is working with 18 arms.
  • 2022-10-18: Two days ago, we noticed some issues in DET2 during the daily quality control checks. After several tests, we concluded that ARM16 is probably broken or loose and today it will be left in simulation. Hence, as of tonight KMOS is working with 19 arms.
  • 2022-10-06: KMOS is working with 20 arms since ARM12 was left in simulation due to loose components at some rotator angles.
  • 2022-10-01: Starting P110, we have introduced a new policy regarding the grading of the OBs in what concerns the number of arms in simulation and the rotator optimisation. The updated information can be found in the Service Mode OB Rules.
  • 2022-09-29: ARM5 showed vignetting on skyflats in some rotator angles suggesting an issue in the arm. Consequently, KMOS is working with 21 arms
  • 2022-09-05: ARM19 showed vignetting on skyflats in some rotator angles suggesting an issue in the arm. Therefore, the arm was left in simulation and hence KMOS is working with 22 arms
  • 2022-08-20: As of today, KMOS is ready for science operations after the cold head maintenance. ARM13 showed vignetting and was left in simulation while ARM19 was recovered. KMOS is working with 23 ARMs.
  • 2022-08-01: IMPORTANT Changes in the calibration plan in what concerns the observation of telluric stars:
    • Starting with Period 111 (Apr - Sep 2023), the calibration plan does not impose a constraint in the airmass of the telluric. The recipe calTrans of the ESO Pipeline can take care of this difference in airmass.
    • From P112 (Oct 2023 - Mar 2024): the calibration plan will not include observations of telluric standard stars for each observation anymore. Instead, users are encouraged to use Molecfit. The observatory will only provide 1 telluric observed in the evening twilight at average airmass.

For some scientific case, the observation of a telluric standard star close in time and airmass to the science may be necessary. In such cases, the execution time for the observation(s) of the telluric standard star (1 per Observing Block of the science target) must be included in the phase 1 proposal and be submitted in concatenation with the science during phase 2.

The Telluric Standard Star Catalogue Search Tool can be used to find appropriate stars (the "Hipparcos/2MASS B stars single" catalogue contains a list of single, B-type, main-sequence stars). The output will provide the magnitude (use I for the IZ grating; J for the YJ grating; H for the H grating; K for the K grating, and “H,K” for the HK grating), spectral type and coordinates needed for the creation of the telluric OB in p2.  

  • 2022-08-01: As of today KMOS is out of operations due to a maintenance of the cooling system. It is expected to be back in operations on Aug 13th.
  • 2022-05-15: ARM19 failed in its movements and was left in simulation. Hence, KMOS is working with 23 ARMs.
  • 2022-03-01: KMOS is fully operational
  • 2022-02-09: KMOS is out of operations due to a problem in one of the cold heads.
  • 2021-10-08: After an intervention, all 24 ARMs are fully functional.
  • 2021-09-10: Important note for the reduced data in the HK grating. The data reduced by pipeline versions 4.0.2 and precedent, have the flux and associated error under-estimated by a factor ~1.6. This bug will be solved in future pipeline and reduced data releases. Please check the header keyword PROCSOFT to know the pipeline version.
  • 2021-07-05: KMOS is currently operating with 17 arms since Arms #2, 14, 15, 18, 20, 22, 24 are in simulation.
  • 2020-08-04 Release of Pipeline Processed and Telluric Corrected KMOS 3D Data Cubes. The KMOS data cubes tagged "KMOS" can be queried and downloaded using the Phase 3 generic query form which provides access to all Phase 3 data, or the Phase 3 spectral data query form with query parameters suited to harvest 3D cube data. KMOS data can be browsed and accessed via the ASP web interface or programmatically.
  • 2019-05-10 KMOS is available for scientific operations with 23 arms as arm #14 is parked and in simulation. The telluric stardard is observed with arms #3, #12 and #18.
  • 2019-05-08 Arms 4 and 8 have been replaced. Unfortunately during the verification activities Arm #14 showed a hard point in the stepping motor of the linear device. It has been parked and is currently in simulation. KMOS can operate with 23 arms till next intervation. Verification activoties are expected to be completed within May 10th 
  • 2019-04-19 KMOS is currently unavailable for science operations. Arm 8 is stuck and deployed in the field of view at its maximum linear extent. An emergincy intervation is foreseen to start on March 27th. 
  • 2018-10-01 Since September 26 Arm #4 started to fail in its movements. The linear motion is irregular showing lateral drifts when the arm is deployed in the FoV. For safety reasons, the arm is put in simulation and KMOS is available for Science Operations with 23 arms until an intervention can be planned.  
  • 2018-08-30 An excraordinary intervation, concluded on Aug 20, was necessary to replace a broken cold head in Spectrograph #3. Following the recommissioning KMOS is avalable for science since Aug 23. However the performances of spectrograph #2 are degraded with astigmatism along the spatial direction (slit width decreased from 13.5 px to 13.2 px) and spectral direction with and degraded resolution of ~10%. The former can be appreciated only under exceptional seeing conditions, when the Nyquist sampling is almost reached. In Specrograph #3 we register a ~10% improve of the spectral resolution
  • 2018-08-09 Due to the failure of a cold head, KMOS experienced a warm up of all the spectrographs and of detector #3. An emergency intervention is already ongoing. As for the current plan, the recomissioning will start on Aug 23.
  • 2018-05-30 The intervations on KMOS and following commissioning have been succesful: Arm 1 and 3 have been replaced and arm 17 fixed. KMOS is now available for Science Operations with all 24 ARMs
  • 2018-05-15 An intervention just started on KMOS to maintain the cold heads and fix/replace the underperforming arms (namely arm #1, #3 and #17). According to the current schedule the recomissioning is foreseen to start on May 23.
  • 2018-01-07 Since December a flux loss has been detected also for ARM #17. The loss depends on the rotator position, at ROT=240 it reaches ~50%. It is a loss of efficiency in the arm but it is calibrated with the internal flats interpolated to the rotator position (standard pipeline procedure).  Therefore, we will leave the arm active until it shall be fixed during the next intervention planned in the second half of May 2018.
  • 2017-11-14 Since October a flux loss up to ~40% happened in ARM #1. Such flux loss depends on the rotator direction and happens at ROT=0 and 15 deg if the instrument is moving from negative positions towards positive ones, and at ROT=-30 and -45 deg when moving in the opposite direction. The origin of the problem is most probably in some loose optical element within the arm itself. The arm will remain anyhow active during the rest of P100 unless it's movement results unreliable. An intervention is planned in P101 to replace both Arm 1 and 3. 
  • 2017-08-08 As a consequence of a power cut, the focus in Spectrograph 3 has changed: the line spread function has a FWHM ~15km/s narrower than before.
  • 2017-07-04 As of P100, a new KARMA version is available (version 2.7). This should be used for the preparation of the KARMA set up files. The new KARMA version allows to allocate 2 arms to the same target, one at the Telescope at Science position and the other at the Telescope at Sky position. This strategy is useful, for example, if the number of targets is less than 24. It has the advantage that the target is observed continuously throughout the whole nodding cycle, thus optimizing the observing time at the telescope. If this strategy is used it should be stated in the Readme file
  • 2017-06-12 In the period May 14-30 a dirt pattern appeared on the focal plane of detector #1. Most probably, it is a flake of paint that detached from the interior of the instrument and fell onto the detector focal plane. The position of the dirt pattern affects a small area and wavelength range in the IFUs #7 and #8, in particular the top left quadrant of IFU #7 and the top right quadrant of #8. 
  • 2017-01-25 No intervention is foreseen to fix/replace Arm #3 in the next period, therefore KMOS will be operative with 23 arms instead of 24 during P100. Arm #4 is used to observe telluric standards for spectrograph 1. 
  • 2017-01-09 Arm 3 is still in simulation and will remain in this state as soon as the instrument can be warmed up and the arm fixed/replaced. Even though the astrometry is indeed re-calibrated, Arm 3 is still showing vignetting on sky with such vignetting being dependent on the arm and rotator position. As before, the arm to take the telluric standard for spectrograph1 is #4. 
  • 2016-09-05 Arm 3 is experiecing again issues: bad precision in prositioning as before plus vignetting. These issues are under investigation and Arm 3 is set to simulation and the arm to take the telluric standard for spectrograph1 is now #4.
  • Since 2016-07-18, arm 3 is showing some problems. It had been set to simulation for a short period (25th July to 4th August)  but is now working again, albeit with a higher uncertainty on the positioning (about 3 pixels).  For that reason, the arm to take the telluric standard for spectrograph1, was changed from #3 to #4.
  • 2015-12-11 Starting from P97, we have included the new keyword NCYCLES in the KMOS_spec_obs_freedither template that can be used to repeat the defined dither cycle including sky observations NCYCLES times. This allows i.e. to stay on a target and take NCYCLES observations, making the design of transit observations easier.
  • 2015-12-10 Starting from P97, darks with DITs corresponding to science obersvations, will no longer be taken by default. The 60s darks, that are used by the pipeline for the bad pixel correction will still be taken on a daily basis. This new procedure allows to remove the constraints on the usable DITs making the design of OBs more flexible and time-efficient.
  • 2015-06-11 A facebook group has been created to exchange ideas, methods, software on KMOS data reduction and analysis.
  • 2015-03-04: After an intervention to upgrade all arms, KMOS was handed back to Science Operation for recommissioning and testing.
  • 2015-01-22 Arm 2 was put in simulation
  • 2014-12-02 Arm 19 was put into simulation
  • 2014-11-21  Arm 8 was put into simulation.
  • 2014-11-18  Arm 4 was put into simulation.
  • 2014-10-06: Arm 5,6,7,and 16 have been replaced, Electronics of Arm4 were changed. All arms except 11 are back in operation.
  • 2014-08-24: Arm 7 in simulation.
  • 2014-08-07: Arm 15 has been recoverd and is back in operation.
  • 2014-08-06: Arm 5 was put in simulation.
  • 2014-05-31: Arm 15 was put in simulation.
  • 2014-05-20: Arms 4 and 11 are temporarily out of operations.
  • 2014-04-01: Arm 11 is back in operation.
  • 2014-02-21: KMOS was successfully recommissioned after an intervention to lower the vibration and to repair the arms that were locked during the last months. Arm 11 is still not available, all other arms work well within the specifications.
  • 2014-01-07: arm11 parked and locked
  • 2013-12-23: arm10 parked and locked
  • 2013-10-15: arm2 parked and locked
  • since 2013-10-01: KMOS officially in operation for P92.
  • 2013-09-21 -- 2013-09-29: Science Verification II
    • arm 16 repaired and working fine
    • arm 13 still parked
    • arm 2 starting to make problems
  • 2013-06-24 -- 2013-06-30: Science Verification I
    • arm 16 was found broken and was blocking part of the field of view.
    • arm 13 was parked due to technical problems
  • 2013-03-11 -- 2013-03-31: Commissioning III
  • 2013-01-17 -- 2013-01-30: Commissioning II
  • 2012-11-21 -- 2012-12-04: Commissioning I